i really need u to solve this problem for me. I am having problems learning how to do this and i need help solving it, if you could send me some instructions on how to break inegualities down and solve the problem i would be very thankful.
M+(-3/4)<= 9/10

Found 2 solutions by stanbon, edjones:
Answer by stanbon(48545)

M+(-3/4)<= 9/10
Add 3/4 to both sides to get:
M <= (9/10)+(3/4)
M <= (18/20) + (15/20)
M <= 33/20
====================
Cheers,
Stan H.

Answer by edjones(7311)

m-(3/4)<= 9/10
40m-30<=36 multiply both sides by 40 to eliminate fractions.
40m-30+30<=36+30
40m<=66
40m/40<=66/40
m<=33/20
Ed